a commitee consists of 10 republicans and 8 democrats. in how many ways can a subcommitee be chosen if it has 5 republicans and 4 democrats
Answer by stanbon(75887)

a commitee consists of 10 republicans and 8 democrats. in how many ways can a subcommitee be chosen if it has 5 republicans and 4 democrats
------
10C5 * 8C4 = 252*70 = 17640
===============================
Cheers,
Stan H.
========================